# Graph Diffusion Transformer for Multi-Conditional Molecular Generation

### v2

@@ Abstract
Inverse molecular design with diffusion models holds great potential for advancements in material and drug discovery. Despite success in unconditional molecule generation, integrating multiple properties such as synthetic score and gas permeability as condition constraints into diffusion models remains unexplored. We present the Graph Diffusion Transformer (Graph DiT) for multi-conditional molecular generation. Graph DiT has a condition encoder to learn the representation of numerical and categorical properties and utilizes a Transformer-based graph denoiser to achieve molecular graph denoising under conditions. Unlike previous graph diffusion models that add noise separately on the atoms and bonds in the forward diffusion process, we propose a graph-dependent noise model for training Graph DiT, designed to accurately estimate graph-related noise in molecules. We extensively validate the Graph DiT for multi-conditional polymer and small molecule generation. Results demonstrate our superiority across metrics from distribution learning to condition control for molecular properties. A polymer inverse design task for gas separation with feedback from domain experts further demonstrates its practical utility.
@@ 1 Introduction
To address these challenges, we project multi-properties into representation by learning, thereby guiding the diffusion process for molecule generation. We propose the Graph Diffusion Transformer (Graph DiT) for graph denoising under conditions. Graph DiT has a condition encoder for property representation learning and a graph denoiser. The condition encoder utilizes a novel clustering-based method for numerical properties and one-hot encoding for categorical ones to learn multi-property representations. The graph denoiser first integrates node and edge features into graph tokens, then denoise these tokens with adaptive layer normalization (AdaLN) in Transformer layers [huang2017arbitrary, peebles2023scalable]. AdaLN replaces the molecular statistics (mean and variance) in each hidden layer with those from the condition representation, effectively outperforming other predictor-based and predictor-free conditioning methods [jo2022score, vignac2022digress, peebles2023scalable], as shown in LABEL:subsec:ablation. We observe that existing forward diffusion processes [vignac2022digress, jo2022score] apply noise separately to atoms and bonds, which may compromise the accuracy of Graph DiT in noise estimation. Hence, we propose a novel graph-dependent noise model that effectively applies noise tailored to the dependencies between atoms and bonds within the graph.
@@ 2.2 Diffusion Model on Graph Data
Diffusion models consist of forward and reverse diffusion processes [ho2020denoising]. We refer to the forward diffusion process as the diffusion process following [ho2020denoising]. The diffusion process $q(G^{1:T}\mid G^{0})=\prod_{t=1}^{T}q(G^{t}\mid G^{t-1})$ corrupts molecular graph data ($G^{0}=G$) into noisy states $G^{t}$. As timesteps $T\rightarrow\infty$, $q(G^{T})$ converges to a stationary distribution $\pi(G)$. The reverse Markov process $p_{\theta}(G^{0:T})=q(G^{T})\prod_{t=1}^{T}p_{\theta}(G^{t-1}\mid G^{t})$ is parameterized by neural networks to gradually denoise the latent states toward the desired data distribution.
@@
3 Multi-Conditional Diffusion Model
@@
3.1 Graph-dependent Noise Model
@@
where $\otimes$ denotes the Kronecker product, $\mathbf{1}_{N}$, $\mathbf{1}_{N}^{\prime}$, and $\mathbf{1}_{N\times N}$ represent the column vector, row vector, and matrix with all 1 elements, respectively. According to Eq. 5, the first $F_{V}$ columns in $\mathbf{Q}_{G}$ determine the node feature transitions based on both node features (first $F_{V}$ rows) and edge features (remaining $N\cdot F_{E}$ rows). Conversely, the remaining $N\cdot F_{E}$ columns determine the edge feature transitions, depending on the entire graph. We introduce a new diffusion noise model applied to $\mathbf{X}_{G}$:
$q(\mathbf{X}_{G}^{t}\mid\mathbf{X}_{G}^{t-1})=\operatorname{Cat}\left(\mathbf{X}_{G}^{t};\mathbf{p}=\mathbf{X}^{t-1}_{G}\mathbf{Q}_{G}^{t}\right).$ (6)
@@
Similar to the definitions of $\mathbf{m}_{V}$ and $\mathbf{m}_{E}$ [vignac2022digress], we leverage the prior knowledge within the training data for the formulation of task-specific matrices, $\mathbf{Q}_{EV}$ and $\mathbf{Q}_{VE}$. Hence, we calculate co-occurrence frequencies of atom and bond types in the training molecular graphs to derive the marginal atom-bond co-occurrence probability distribution. For each bond type, each row in $\mathbf{m}_{EV}$ represents the probability of co-occurring atom types. $\mathbf{m}_{VE}$ is the transpose of $\mathbf{m}_{EV}$ and has a similar meaning. Subsequently, we define $\mathbf{Q}_{EV}=\bar{\alpha}^{t}\mathbf{I}+(1-\bar{\alpha}^{t})\mathbf{1}\mathbf{m}_{EV}^{\prime}$ and $\mathbf{Q}_{VE}=\bar{\alpha}^{t}\mathbf{I}+(1-\bar{\alpha}^{t})\mathbf{1}\mathbf{m}_{VE}^{\prime}$.
3.2 Denoising Generation with Graph Diffusion Transformer
@@
where $s$ denotes the scale of conditional guidance. Unlike classifier-free guidance [ho2022classifier], which typically predicts noise, we directly estimate $p_{\theta}(\tilde{G}^{0}\mid G^{t},\mathcal{C})$. We utilize the Graph DiT, $f_{\theta}(G^{t},\mathcal{C})$, for both $p_{\theta}(\tilde{G}^{0}\mid G^{t})$ and $p_{\theta}(\tilde{G}^{0}\mid G^{t},\mathcal{C})$. Here, $f_{\theta}(G^{t},\mathcal{C}=\emptyset)$ computes the unconditional probability, substituting original embeddings for a condition with a learnable dropping embedding. To learn this dropping embedding, we randomly replace elements in $\mathcal{C}$ with a special dropping value during training. Graph DiT $f_{\theta}(G^{t}=\mathbf{X}_{G}^{t},\mathcal{C})$ comprises two components: the condition encoder and the graph denoiser. An overview of the architecture is presented in Figure 2.
@@
Given the noisy graph at timestep $t$, the graph tokens are first encoded into the hidden space as $\mathbf{H}=\operatorname{Linear}(\mathbf{X}_{G}^{t})$, where $\mathbf{H}\in\mathbb{R}^{N\times D}$. We then adapt the standard Transformer layers [vaswani2017attention] with self-attention and multi-layer perceptrons (MLP), but replace the layer normalization with adaptive ones controlled by the conditional representation. This adaptive layer normalization is denoted by $\operatorname{AdaLN}$ [huang2017arbitrary, peebles2023scalable]: $\mathbf{H}=\operatorname{AdaLN}(\mathbf{H},\mathbf{c}$). For each row $\mathbf{h}$ in $\mathbf{H}$:
@@ 4.1 Experimental Setup
Table 1: Generation of 10K Polymers: Results on a numerical synthesizability score (Synth.) and three numerical properties (gas permeability for O2, N2, CO2). MAE is calculated between input conditions and generated properties. Best results are highlighted.

### v3

# Graph Diffusion Transformers for Multi-Conditional Molecular Generation

@@ Abstract
Inverse molecular design with diffusion models holds great potential for advancements in material and drug discovery. Despite success in unconditional molecular generation, integrating multiple properties such as synthetic score and gas permeability as condition constraints into diffusion models remains unexplored. We present the Graph Diffusion Transformer (Graph DiT) for multi-conditional molecular generation. Graph DiT integrates an encoder to learn numerical and categorical property representations with the Transformer-based denoiser. Unlike previous graph diffusion models that add noise separately on the atoms and bonds in the forward diffusion process, Graph DiT is trained with a novel graph-dependent noise model for accurate estimation of graph-related noise in molecules. We extensively validate Graph DiT for multi-conditional polymer and small molecule generation. Results demonstrate the superiority of Graph DiT across nine metrics from distribution learning to condition control for molecular properties. A polymer inverse design task for gas separation with feedback from domain experts further demonstrates its practical utility.
@@ 1 Introduction
To address these challenges, we project multi-properties into representations by learning, thereby guiding the diffusion process for molecule generation. We propose the Graph Diffusion Transformer (Graph DiT) for graph denoising under conditions. Graph DiT has a condition encoder for property representation learning and a graph denoiser. The condition encoder utilizes a novel clustering-based method for numerical properties and one-hot encoding for categorical ones to learn multi-property representations. The graph denoiser first integrates node and edge features into graph tokens, then denoise these tokens with adaptive layer normalization (AdaLN) in Transformer layers [huang2017arbitrary, peebles2023scalable]. AdaLN replaces the molecular statistics (mean and variance) in each hidden layer with those from the condition representation, effectively outperforming other predictor-based and predictor-free conditioning methods [jo2022score, vignac2022digress, peebles2023scalable], as shown in LABEL:subsec:ablation. We observe that existing forward diffusion processes [vignac2022digress, jo2022score] apply noise separately to atoms and bonds, which may compromise the accuracy of Graph DiT in noise estimation. Hence, we propose a novel graph-dependent noise model that effectively applies noise tailored to the dependencies between atoms and bonds within the graph.
@@ 2.2 Diffusion Model on Graph Data
Diffusion models consist of forward and reverse diffusion processes [ho2020denoising]. We refer to the forward diffusion process as the diffusion process following [ho2020denoising]. The diffusion process $q(G^{1:T}\mid G^{0})=\prod_{t=1}^{T}q(G^{t}\mid G^{t-1})$ corrupts molecular graph data ($G^{0}=G$) into noisy states $G^{t}$. As timesteps $T\rightarrow\infty$, $q(G^{T})$ converges a stationary distribution $\pi(G)$. The reverse Markov process $p_{\theta}(G^{0:T})=q(G^{T})\prod_{t=1}^{T}p_{\theta}(G^{t-1}\mid G^{t})$, parameterized by neural networks, gradually denoises the latent states toward the desired data distribution.
@@
3 Multi-Conditional Graph Diffusion Transformers
@@
3.1 Graph-Dependent Noise Models
@@
where $\otimes$ denotes the Kronecker product, $\mathbf{1}_{N}$, $\mathbf{1}_{N}^{\prime}$, and $\mathbf{1}_{N\times N}$ represent the column vector, row vector, and matrix with all 1 elements, respectively. According to Eq. 5, the first $F_{V}$ columns in $\mathbf{Q}_{G}$ determine the node feature transitions based on both node features (first $F_{V}$ rows) and edge features (remaining $N\cdot F_{E}$ rows). Conversely, the remaining $N\cdot F_{E}$ columns determine the edge feature transitions, depending on the entire graph. We introduce a new diffusion noise model:
$q(\mathbf{X}_{G}^{t}\mid\mathbf{X}_{G}^{t-1})=\widetilde{\operatorname{Cat}}\left(\mathbf{X}_{G}^{t};\tilde{\mathbf{p}}=\mathbf{X}^{t-1}_{G}\mathbf{Q}_{G}^{t}\right),$ (6)
where $\tilde{\mathbf{p}}$ is the unnormalized probability and $\widetilde{\operatorname{Cat}}$ denotes categorical sampling: The first $F_{V}$ columns of $\tilde{\mathbf{p}}$ are normalized to sample $\mathbf{X}_{V}^{t}$, while the remaining $N\cdot E$ dimensions are reshaped and normalized to sample edges $\mathbf{X}_{E}^{t}$. These components are combined to form $\mathbf{X}_{G}^{t}$, completing the $\widetilde{\operatorname{Cat}}$ sampling.
@@
Similar to the definitions of $\mathbf{m}_{V}$ and $\mathbf{m}_{E}$ [vignac2022digress], we leverage the prior knowledge within the training data for the formulation of task-specific matrices, $\mathbf{Q}_{EV}$ and $\mathbf{Q}_{VE}$. We calculate co-occurrence frequencies of atom and bond types in training molecular graphs to obtain the marginal atom-bond co-occurrence probability distribution. For each bond type, each row in $\mathbf{m}_{EV}$ represents the probability of co-occurring atom types. $\mathbf{m}_{VE}$ is the transpose of $\mathbf{m}_{EV}$ and has a similar meaning. Subsequently, we define $\mathbf{Q}_{EV}=\bar{\alpha}^{t}\mathbf{I}+(1-\bar{\alpha}^{t})\mathbf{1}\mathbf{m}_{EV}^{\prime}$ and $\mathbf{Q}_{VE}=\bar{\alpha}^{t}\mathbf{I}+(1-\bar{\alpha}^{t})\mathbf{1}\mathbf{m}_{VE}^{\prime}$.
3.2 Denoising Models with Multi-Property Conditions
@@
where $s$ denotes the scale of conditional guidance. Unlike classifier-free guidance [ho2022classifier], which typically predicts noise, we directly estimate $p_{\theta}(\tilde{G}^{0}\mid G^{t},\mathcal{C})$. We one one denoising model $f_{\theta}(G^{t},\mathcal{C})$ for both $p_{\theta}(\tilde{G}^{0}\mid G^{t})$ and $p_{\theta}(\tilde{G}^{0}\mid G^{t},\mathcal{C})$. Here, $f_{\theta}(G^{t},\mathcal{C}=\emptyset)$ computes the unconditional probability by substituting the original conditional embeddings with the null value. During training, we randomly drop the condition with a ratio, i.e., $\mathcal{C}=\emptyset$, to learn the embedding of the null value. $f_{\theta}(G^{t}=\mathbf{X}_{G}^{t},\mathcal{C})$ comprises two components: the condition encoder and the graph denoiser. An overview of the architecture is presented in Figure 2.
@@
Given the noisy graph at timestep $t$, the graph tokens are first encoded into the hidden space as $\mathbf{H}=\operatorname{Linear}(\mathbf{X}_{G}^{t})$, where $\mathbf{H}\in\mathbb{R}^{N\times D}$. We then adapt the standard Transformer layers [vaswani2017attention] with self-attention and multi-layer perceptrons (MLP), but replace the normalization with the adaptive layer normalization ($\operatorname{AdaLN}$) controlled by the representations of the conditions [huang2017arbitrary, peebles2023scalable]: $\mathbf{H}=\operatorname{AdaLN}(\mathbf{H},\mathbf{c}$). For each row $\mathbf{h}$ in $\mathbf{H}$:
@@ 4.1 Experimental Setup
Table 1: Multi-Conditional Generation of 10K Polymers: Results on the synthetic score (Synth.) and three numerical properties (gas permeability for O2, N2, CO2). MAE is calculated between the input conditions and the properties of the generated polymers using Oracles. Best results are highlighted.